Last year Kruse Corp had $410,000 of assets (which is equal to its total invested capital), $403,000 of sales, $28,250 of net income, and a debt-to-total-capital ratio of 39%. The new CFO believes the firm has excessive fixed assets and inventory that could be sold, enabling it to reduce its total assets and total invested capital to $252,500. The firm finances using only debt and common equity. Sales, costs, and net income would not be affected, and the firm would maintain the same capital structure (but with less total debt). By how much would the reduction in assets improve the ROE? Do not round your intermediate calculations.
Answer #1

Answer:

ROE before reduction in assets:

Total assets = $410,000

Debt to total capital ratio = 39%

Equity to total capital ratio = 1 - 39% = 61%

Equity = 410000 * 61% = $250,100

Net Income = $28,250

ROE = Net Income / Equity = 28250 / 250100 = 11.2955%

After reduction in assets:

Total assets = $252,500

Net Income is not affected and is same at = $28,250

Capital structure is same.

New Equity = 252500 * 61% = $154,025

New ROE = 28250 / 154025 = 18.3412%

Improvement in ROE = 18.3412% - 11.2955% = 7.05%

Total Assets Equity (61%) Net Income ROE $410,000 | Before reduction in assets | $250,100 | $ % 28,250 11.2955 After reductio